# Linying Lv

### v2

@@ 3.2.3 What Content in Analyst Reports is Valued by Market Investors?
The Shapley value decomposition, as illustrated in Figure 2, reveals a striking hierarchy of topic importance in the written reports. The most valuable topics are Income Statement Analyses and Financial Ratio Analyses, each contributing to 67% and 45% of text-based $R^{2}_{\mathrm{OOS}}$. Following these, Investment Thesis and Valuation constitute the second tier of importance, with Shapley values of 1.76% and 1.51%, respectively. The remaining topics exhibit minimal or even negative Shapley values, suggesting negligible explanatory power for contemporaneous market reactions to analyst reports.
@@
The robustness of these findings is further reinforced by additional analyses. As demonstrated in Figure A4, scaling Shapley values by sentence count and token length does not alter the preeminence of Income Statement Analyses. Figure A5 reveals consistent topic rankings across industries.
Figure A6 in the Online Appendix shows how the text-based $R^{2}_{\mathrm{OOS}}$ changes over time. A significant drop in $R^{2}_{\mathrm{OOS}}$ is observed between 2020 and 2021, likely due to the disruptive effects of the COVID-19 pandemic on financial markets and the broader economy. Despite the volatility, the scaled Shapley value for the Income Statement category, represented by the red line, remains relatively stable. Analyst discussions of income statements consistently provide valuable information, even during market disruptions.
@@
The relative importance of topics could stem from analysts’ expertise in certain areas (supply effect) or from investors’ varying responses to different topics (demand effect). Figure A7 compares Shapley values across analysts of different experience levels and brokerage firm sizes. The consistency in topic importance across all four subsamples (experienced vs. inexperienced analysts, large vs. small brokerages) suggests that topic rankings are not driven by differences in analysts’ topical expertise.
Building upon the initial findings, I proceed to study the interpretation role and discovery role of analysts in analyzing income statements. Drawing inspiration from Huang et al. (2018) and Blankespoor et al. (2020), I distinguish between the analyst’s role in collecting and generating otherwise unavailable information (discovery role) and their ability to integrate and process available information into more interpretable signals (interpretation role).
@@
To facilitate the analysis, I categorize Income Statement sentences based on two criteria: (1) Information Acquisition versus Information Interpretation, and (2) Realized Income versus Expected Income. This categorization is achieved through a prompt-based classification system, as detailed in Prompt 3.
@@
Figure 3 provides a detailed breakdown of $R^{2}_{\mathrm{OOS}}$ across 18 topics, with Income Statement subcategories. Income interpretation contributes three times as much to $R^{2}_{\mathrm{OOS}}$ as income acquisition, suggesting that markets value context and explanation over raw data. The contribution from realized income discussions is triple that of income expectations, indicating markets place greater value on analysis of actual results than on forward-looking projections. These patterns demonstrate that markets particularly value analysts’ interpretation of realized income, more than their reporting of public information or future expectations.
This insight is reinforced by the findings in section 3.2.2, which show that analyst reports issued promptly after earnings announcements are the most informative. Taken together, these results suggest that the market places the highest value on analysts who can quickly digest financial information from earnings announcements and provide insightful interpretations.
@@ 3.2.4 Comparison of Revision Reports and Reiteration Reports
Researchers have documented a robust relationship between market prices and analysts’ forecast revisions (e.g., Gleason and Lee, 2003; Asquith et al., 2005; Bradshaw et al., 2021). Analysts’ forecast revisions provide a significant and timely source of information to the financial market and may constitute a more valuable signal than reports that merely reiterate previous forecasts.
Therefore, I compare the information content of analyst report text associated with forecast revisions and forecast reiterations. Following the criteria of Huang et al. (2014), I consider a report as a revision report if its release date is within two days of the I/B/E/S forecast announcement date. For each forecast target in recommendations, target prices, and earnings forecasts, I categorize the report independently. Consequently, a single report may be classified as both an earnings forecast revision report and a recommendation reiteration report.
Table 5 summarizes the out-of-sample performance of models with numerical and textual inputs using both reiteration and revision samples. Recommendation revisions are rare events compared to target price revisions and earnings forecast revisions. Earnings forecast revisions are the most frequent, comprising 56.57% of the reports. In contrast, recommendation revision reports and target price revision reports constitute only 2.29% and 31.10%, respectively. When stock recommendations are revised, the explanatory power of various information sources becomes particularly pronounced. The quantitative forecast revision measures alone achieve an $R^{2}_{\mathrm{OOS}}$ of 14.99%. The textual content of analyst reports accompanying recommendation revisions explains a larger portion of CAR variation, with an $R^{2}_{\mathrm{OOS}}$ of 16.80%. The combination of both quantitative and qualitative information further enhances the model’s explanatory power, resulting in a $R^{2}_{\mathrm{OOS}}$ of 22.63%. This finding complements the literature that attributes large stock price changes to analysts’ recommendation revisions (Loh and Stulz, 2011; Bradley et al., 2014).
Target price revisions are considered incrementally informative due to the discreteness and potential biases in analysts’ recommendations (Bradshaw, 2002; Brav and Lehavy, 2003). While reports with target price reiterations yield low information content (combined $R^{2}_{\mathrm{OOS}}$ of 3.58% for three-day CAR), revision reports demonstrate substantial value ($R^{2}_{\mathrm{OOS}}$ = 20.88%, t-stat = 11.18).
As Kothari et al. (2016) note, analysts’ earnings forecast revisions convey both public and private information about cash flow news in financial markets and are the most extensively studied type of analyst output. For models with text from earnings forecast reiteration reports as input, the $R^{2}_{\mathrm{OOS}}$ is insignificant. When focusing on analyst earnings forecast revisions, the $R^{2}_{\mathrm{OOS}}$ for both numerical and textual inputs becomes significant. The combination yields an $R^{2}_{\mathrm{OOS}}$ of 16.37%, with a t-statistic of 11.42, underscoring the greater value of revisions over reiterations.
Collectively, Table 5 shows that report content is more informative when it accompanies forecast revisions rather than reiterations. Additionally, analyst reports associated with recommendation revisions demonstrate the highest value.
@@ 3.2.5 Comparison with Analyst Report Sentiment
Huang et al. (2014) emphasize the information content of analyst report text by extracting tone using a Naive Bayes machine learning approach. I compare report-level tone measures derived from both Naive Bayes and BERT classification. Following Huang et al. (2014), I manually label 10,000 sentences as training data for both models. Each sentence is categorized as positive, neutral, or negative using these pre-trained models, and I average the tone at the report level to create $Tone_{N}B$ and $Tone_{B}ERT$ measures. I also construct tone measures separately for Income Statement topics and non-Income Statement topics.
@@
Table 6 presents a comparison of performance between text tone measures and text embeddings. Panel A summarizes the OLS regression of three-day $CAR$ on the tone measures. The coefficient estimate for $Tone\_NB$ in column (1) is 0.019, comparable to the 0.0208 reported by Huang et al. (2014). A one standard deviation increase in the favorableness of the text tone results in an additional $CAR$ of 49 basis points, comparable with the 41 basis points increase in two-day abnormal return reported in Huang et al. (2014). Column (2) summarizes the results of regressing $CAR$ on tone measures extracted within Income Statement topics and non-Income Statement topics. Both coefficients of tone are significantly positive. Regression results with firm-fixed and year-fixed effects are consistent, as reported in column (3).
Columns (4)-(6) summarizes regression results with BERT-based tone measures as explanatory variables. The coefficient estimates’ magnitude increases dramatically, along with significance levels. Economically, a one standard deviation increase in BERT-based text tone favorableness results in an additional three-day $CAR$ of 86 basis points. When both tone measures are included in an OLS regression in column (7), coefficient estimates of Naive Bayes-based tone measure become insignificant, suggesting that BERT model captures sentiment more accurately and effectively.
Panel B of Table 6 compares the $R^{2}_{\mathrm{OOS}}$ of tone measures and embeddings. The $R^{2}_{\mathrm{OOS}}$ of tone inputs ranges from 0.05% to 3.68% in the first four rows. The $R^{2}_{\mathrm{OOS}}$ improves to 9.49% and 12.27% after incorporating forecast revisions and text embeddings. The bottom two rows of Panel B report comparisons of $R^{2}_{\mathrm{OOS}}$ and DM tests between models with “Tone + Rev + text” versus “Rev + text.” The difference between the $R^{2}_{\mathrm{OOS}}$ of models with and without tone measures is not statistically distinguishable from zero.
The empirical results in Table 6 highlight two key advantages of LLMs over conventional NLP methods. First, tone measures extracted using LLMs better explain contemporaneous stock returns than those from the Naive Bayes algorithm. Second, neither BERT nor Naive Bayes tone measures add incremental value beyond text embeddings, suggesting that LLMs capture richer contextual information than simple tone metrics.
3.2.6 Information Content with Alternative Machine Learning Models
In this Section, I examine the information content using three alternative machine learning algorithms, applying the same processes as in Ridge regressions. Each model uses full-context report embeddings as inputs, with the target variable being the three-day $CAR$.
@@ 3.2.6 Information Content with Alternative Machine Learning Models
Table 7 presents results from PLS regressions. PLS effectively extracts common components that correlate with the target variable. If text embeddings contain strong common components that predict stock returns, PLS regression should capture this information. The PLS regressions achieve an $R^{2}_{\mathrm{OOS}}$ of 8.26%, with year-to-year variations similar to those in Ridge models. DM tests confirm that the $R^{2}_{\mathrm{OOS}}$ is significant at the 1% level across all years.
Next, I replicate the analyses with the XGBoost algorithm. Gradient boosting combines weak estimations to form a strong estimation by constructing sequential trees. Each tree is responsible for training a small sample to fit the residuals of the previous trees and updating the model by learning new information. Therefore, if the high-dimensional feature spaces of text embeddings contain segmented information with discrete importance, XGBoost can potentially identify the most informative aspects of the analyst report embeddings for explaining contemporaneous stock returns. Compared to Ridge and PLS models, XGBoost is capable of capturing complex nonlinear relationships between analyst report embeddings and the target variable. As summarized in Table 7, the XGBoost algorithm underperforms Ridge regressions with an overall $R^{2}_{\mathrm{OOS}}$ of 5.97% (DM t-stat = 6.76). The relatively poor performance could be attributed to the inherent complexity of language, where information is distributed across the embedding space rather than concentrated in specific dimensions.
I then proceed with Neural Network models, which form the core of LLMs and the transformer architecture. In a Neural Network, nodes are interconnected, mimicking the neuron structure of the human brain. Since embeddings are encoded as nodes within the Neural Network layers, I expect this architecture to capture long-range relationships between words and the complex contextual information embedded in the text. I construct the Neural Network estimations as an ensemble of five models for each NN configuration. As shown in Table 7, the $R^{2}_{\mathrm{OOS}}$ of Neural Network models ranges from 10.42% to 12.12%, achieving the highest performance among all tested ML algorithms. Performance tends to improve with network depth, aligning with the virtue of complexity principle proposed by Kelly et al. (2022).
The experiments with machine learning models yield noteworthy findings. First, the information content of report text, as indicated by $R^{2}_{\mathrm{OOS}}$, achieves around 10% in certain models, which is significant considering the low signal-to-noise ratio in stock returns. Second, the superior performance of deep learning architectures emphasizes the importance of contextual information and word interdependencies in textual analyses.
@@ 3.3 Information Value of Analyst Reports
How to translate statistical measure $R^{2}_{\mathrm{OOS}}$ into economic terms? In this section, I quantify the information value of analyst reports using the measure introduced in Section 2.3. This measure reflects the profit derived from a reduction in price uncertainty due to an informative signal, while accounting for price impact. In other words, both greater information content in analyst reports and lower price impact will lead to higher estimated information value.
@@
I first summarize the value of information from analyst reports in Table 8, covering the period from 2015 to 2023. Both the mean and standard deviation are derived using the delta method. For S&P 100 constituent stocks, the average three-day window information value from sell-side analysts is \$0.47 million, with a standard error of \$0.05 million. The mean information value of text alone is \$0.38 million, and the mean value for revisions is \$0.34 million. Considering that a stock typically has reports released on an average of 15 days each year, the annualized lower bound investors would expect to profit from early access to analyst reports on an S&P 100 firm is \$6.89 million. The narrow standard errors and confidence intervals across different information values indicate consistent estimates. The average price impact ($\widehat{\lambda}_{it}/P_{it_{-}}$) is 0.34, indicating that \$1 million dollar order pushes the stock price up by 0.034%. These statistics illustrate that information from analysts is economically meaningful, with both textual information and revisions contributing significantly.
Figure 4 describes the evolution of information value from analyst reports over time. The solid line represents the strategic value of combined numerical and textual information, while the dashed line shows the value of textual information alone. Both lines exhibit significant fluctuations over the period. Despite these fluctuations, there is a general upward trend in information value. For most of the period, the combined value is higher than the textual value alone, suggesting that numerical information often complements the textual content in analyst reports.
The shaded area represents the 95% confidence interval. After 2020, there appears to be greater variability in the information value, as indicated by the wider confidence intervals and more pronounced peaks and troughs. This period corresponds with the global COVID-19 pandemic, which likely introduced additional uncertainty and volatility in financial markets, impacting the value of information from analyst reports.
These patterns indicate that the information value of analyst reports is dynamic, and influenced by market conditions. The relationship between combined and text-only information value fluctuates over time, highlighting the importance of considering both numerical and textual components when analyzing analyst reports.
@@
In the Online Appendix, Table A7 provides the summary statistics of alternative estimates for information value. Using high-frequency measures of realized return volatility and different algorithms for signing trades as buys or sells, the estimates for information value from analyst reports range from \$0.42 million to \$0.64 million. Figure A8 illustrates the time trends of these three alternative estimates, revealing consistent patterns in the evolution of information value over time.
@@ 3.3.1 The Value of Analyst Information in the Cross-section of Stocks
The information value of sell-side analysts likely vary in the cross-section of stocks. Specifically, the costs of information processing and dissemination may differ by firm size (Zeghal, 1984). For example, analyst opinions on smaller stocks may trigger larger market reactions, as these stocks generally have less publicly available information (Stickel, 1995). I plot the scatter of information value and market capitalization by stock in Figure A9. There is a general upward trend, indicating that stocks with higher market equity tend to have higher information values. However, the relationship is not strictly linear, as there are notable exceptions. I formally test the relationship by regressing the value of information on stock size, firm fixed effects, and year fixed effects. I also control for the book-to-market ratio in the regressions.
Table 9 shows the results. Stocks with larger market equity have significantly higher information value from analysts. Economically, results in column (4) indicate a 1% increase in market equity is associated with a 0.864% increase in analyst information value. Columns (5)-(8) reveal that the information value of report text is more sensitive to size compared to the combined numerical and textual information value. I further investigate whether the positive size elasticity of information value is driven by higher explained return volatility or lower price impact of large stocks. The decomposition can be expressed logarithmically as:
@@
Panel B of Table 9 shows that the higher information value for large stocks stems primarily from their greater liquidity. Analyst reports of large stocks do not explain higher contemporaneous return volatility, yet their value increases with stock size due to enhanced liquidity.

### v3

@@ 3.2.3 What Content in Analyst Reports is Valued?
Figure 2 illustrates the Shapley value decomposition, revealing a clear hierarchy in topic importance. Income Statement Analyses and Financial Ratio Analyses emerge as the most valuable components, contributing 67% and 45% of text-based $R^{2}_{\mathrm{OOS}}$, respectively. Investment Thesis and Valuation form a second tier of importance, with Shapley values of 1.76% and 1.51%. The remaining topics show minimal or negative Shapley values, indicating limited market reactions.
@@
The findings’ robustness is supported by several additional analyses. Figure A4 shows that normalizing Shapley values by sentence count and token length maintains Income Statement Analyses’ dominance. Figure A5 demonstrates consistency in topic rankings across different industries.
The evolution of text-based $R^{2}_{\mathrm{OOS}}$ is depicted in Figure A6 of the Online Appendix. While a notable decline in $R^{2}_{\mathrm{OOS}}$ occurs between 2020 and 2021, likely reflecting COVID-19 pandemic disruptions in financial markets, the scaled Shapley value for Income Statement analysis (shown by the red line) exhibits remarkable stability. This persistence suggests that analyst commentary on income statements maintains its informational value even during market turbulence.
@@
To disentangle whether topic importance stems from analysts’ expertise in certain areas (supply effect) or investors’ response (demand effect), Figure A7 compares Shapley values across analyst experience levels and brokerage firm sizes. The consistent topic hierarchy across all four subsamples (experienced/inexperienced analysts and large/small brokerages) indicates that topic rankings are not driven by variations in analyst expertise.
@@
I proceed to examine analysts’ dual roles in income statement analysis: their discovery role in collecting and generating otherwise unavailable information and their interpretation role in integrating and processing available information into more interpretable signals (Huang et al., 2018; Blankespoor et al., 2020). To facilitate the analysis, I categorize Income Statement sentences based on: (1) Information Acquisition versus Information Interpretation, and (2) Realized Income versus Expected Income. This categorization is achieved through a prompt-based classification system, as detailed in Prompt 3.
@@
Figure 3 provides a granular breakdown of $R^{2}_{\mathrm{OOS}}$ across 18 topics, including Income Statement subcategories. Income interpretation contributes three times as much to $R^{2}_{\mathrm{OOS}}$ as income acquisition, suggesting that markets value context and explanation over raw figures. The contribution from realized income discussions is triple that of income expectations, indicating markets place greater value on actual results than forward-looking projections. These patterns emphasizes that markets particularly value analysts’ ability to interpret realized income, beyond mere information reporting or future expectations.
These findings align with the results from section 3.2.2, where analyst reports released immediately after earnings announcements demonstrate the highest information content. Taken together, these results suggest that the market places the highest value on analysts who can rapidly digest earnings announcement information and provide insightful interpretations.
@@ 3.2.4 Comparison of Revision Reports and Reiteration Reports
Given the established relationship between market prices and analyst forecast revisions (e.g., Gleason and Lee, 2003; Asquith et al., 2005; Bradshaw et al., 2021), I examine the differential information content between revision and reiteration reports. Following the criteria of Huang et al. (2014), I classify reports as revisions when released within two days of the I/B/E/S forecast announcement date. This classification is applied independently across three forecast targets—recommendations, target prices, and earnings. Therefore, a single report may be classified as both an earnings forecast revision report and a recommendation reiteration report.
Table 6 presents out-of-sample model performance across reiteration and revision samples. The frequency of revisions varies considerably: earnings forecast revisions are most common (56.57% of reports), followed by target price revisions (31.10%), while recommendation revisions are relatively rare (2.29%). Recommendation revisions exhibit particularly strong explanatory power. Quantitative forecast revision measures achieve an $R^{2}_{\mathrm{OOS}}$ of 14.99% when accompanying recommendation revisions, while report text yields 16.80%. The combination enhances explanatory power to 22.63%. This finding complements the literature that attributes large stock price changes to analysts’ recommendation revisions (Loh and Stulz, 2011; Bradley et al., 2014).
Target price revisions, noted for their incremental information value due to the limitations of discrete recommendations (Bradshaw, 2002; Brav and Lehavy, 2003), show marked differences between revisions and reiterations. While reiteration reports demonstrate minimal information content (combined $R^{2}_{\mathrm{OOS}}$ of 3.58%), revision reports exhibit substantial value ($R^{2}_{\mathrm{OOS}}$ = 20.88%, t-stat = 11.18).
Earnings forecast revisions show a similar pattern. Reiteration reports yield insignificant $R^{2}_{\mathrm{OOS}}$, while revision reports demonstrate significant explanatory power. The combination of numerical and textual inputs for revision reports achieves an $R^{2}_{\mathrm{OOS}}$ of 16.37% (t-statistic = 11.42), highlighting revisions’ superior value over reiterations.
Overall, Table 6 shows that report content carries greater information content when accompanying forecast revisions rather than reiterations, with recommendation revision reports exhibiting the highest value.
@@ 3.2.5 Comparison with Analyst Report Sentiment
The sentiment analysis serves two objectives: to evaluate how LLMs perform compared to traditional NLP methods, and to assess the relative information content of embeddings versus tone measures. I compare tone measures derived from both Huang et al. (2014)’s Naive Bayes approach and BERT classification. I manually label 10,000 sentences as training data for both models. After training the models to categorize sentences as positive, neutral, or negative, I average these classifications at the report level to create $Tone\_NB$ and $Tone\_BERT$ measures.
@@
Table 7 contrasts the performance of text tone measures against text embeddings. Panel A presents OLS regression results of three-day $CAR$ on tone measures. Column (1) shows a coefficient estimate of 0.022 for $Tone\_NB$, aligning with Huang et al. (2014)’s reported 0.0208. A one standard deviation increase in text tone favorableness corresponds to a 54 basis point increase in $CAR$, comparable to their reported 41 basis point increase in two-day abnormal returns. Column (2) reveals significantly positive coefficients for tone measures for both Income Statement and non-Income Statement topics. The results remain consistent under firm-fixed and year-fixed effects in column (3). BERT-based tone measures (columns 4-6) demonstrate substantially larger coefficient estimates with higher significance levels. A one standard deviation increase in BERT-based tone favorableness yields a 95 basis point increase in three-day $CAR$. When both measures are included (column 7), Naive Bayes-based tone coefficients become insignificant, indicating BERT’s superior sentiment capture.
Panel B of Table 7 compares the $R^{2}_{\mathrm{OOS}}$ across models. Tone inputs alone achieve $R^{2}_{\mathrm{OOS}}$ ranging from 0.05% to 3.68% in the first four rows. Including forecast revisions and text embeddings improves $R^{2}_{\mathrm{OOS}}$ to 9.49% and 12.27%, respectively. The panel’s final rows show that adding tone measures to text embedding models yields no statistically significant improvement in $R^{2}_{\mathrm{OOS}}$.
The results highlight two key advantages of LLMs over conventional NLP methods: stronger explanatory power for stock returns through better sentiment capture, and the ability to extract richer textual information beyond simple sentiment measures.
3.2.6 Alternative Machine Learning Models
I examine the information content using three alternative machine learning algorithms, following the same procedure of Ridge regressions. Results are presented in Table 8.
@@ 3.2.6 Alternative Machine Learning Models
PLS regressions extract common components that correlate with the target variable. If text embeddings contain strong common components that correlates with stock returns, PLS should outperform. Using this method, I achieve an $R^{2}_{\mathrm{OOS}}$ of 8.26%, with year-to-year variations similar to Ridge models. DM tests confirm significance at the 1% level across all years.
The XGBoost algorithm, which builds sequential trees to combine weak estimations into stronger ones, allows for testing whether text embeddings contain discrete informational components of varying importance. Despite its ability to capture nonlinear relationships, XGBoost underperforms Ridge regressions with an $R^{2}_{\mathrm{OOS}}$ of 5.97% (DM t-stat = 6.76). This relatively weaker performance suggests that linguistic information is distributed across the embedding space rather than concentrated in specific dimensions.
Neural Network models, which underpin LLMs and transformer architectures, show superior performance. Given their interconnected node structure mirroring human neural networks, these models are well-suited to capture word relationships and contextual information. Using ensembles of five models for each configuration, Neural Networks achieve the highest $R^{2}_{\mathrm{OOS}}$ among tested algorithms, ranging from 10.42% to 12.12%. Performance improves with network depth, consistent with Kelly et al. (2022)’s virtue of complexity principle.
Two key findings emerge from these experiments. First, the significant $R^{2}_{\mathrm{OOS}}$ across models demonstrates meaningful information content in report text, particularly given the low signal-to-noise ratio in stock returns. Second, although Neural Networks achieve the highest performance, most of the textual information can be captured by simple linear models like Ridge regressions.
@@ 3.3 Information Value of Analyst Reports
To translate the statistical measure $R^{2}_{\mathrm{OOS}}$ into economic terms, this section quantifies the information value of analyst reports using the methodology established in Section 2.3. This measure captures the economic profit generated from reduced price uncertainty, while incorporating price impact considerations. The measure’s construction implies that both higher information content and lower price impact contribute positively to the estimated information value.
@@
Table 9 presents the statistical evidence estimated with the delta method. For S&P 100 constituent stocks, sell-side analysts generate an average three-day window information value of \$0.47 million (SE = \$0.05 million). This value can be decomposed into its constituent components: textual content contributes \$0.38 million, while numerical revisions account for \$0.34 million. Extrapolating from the typical frequency of analyst coverage—approximately 15 report days per year—the annualized lower bound investors would expect to profit from early access to analyst reports on an S&P 100 firm amounts to \$6.89 million. The precision of these estimates is supported by consistently narrow standard errors and confidence intervals across different information value measures. The average price impact ($\widehat{\lambda}_{it}/P_{it_{-}}$) is 0.34, suggesting that a \$1 million order pushes the stock price up by 0.034%. The statistics demonstrate that analyst information value is economically meaningful, with both textual content and numerical forecasts contributing significantly.
Figure 4 illustrates the dynamics of analyst report information value. The analysis tracks both the strategic value of combined numerical and textual information (solid line) and the isolated value of textual information (dashed line). While both measures exhibit substantial temporal variation, they demonstrate an overall upward trend. The consistently higher combined value relative to textual value suggests complementarity between numerical and textual components.
The 95% confidence interval reveals increased variability in information value post-2020, characterized by wider confidence bounds and more pronounced peaks and troughs. The heightened volatility coincides with the COVID-19 pandemic period, which introduced substantial market uncertainty. The patterns demonstrate that the dynamic nature of analyst information value and its sensitivity to market conditions.
@@
The Online Appendix provides additional robustness checks. Table A7 presents alternative information value estimates using high-frequency realized return volatility measures and varied trade-signing algorithms, yielding estimates ranging from \$0.42 million to \$0.64 million. Figure A8 demonstrates that these alternative specifications produce consistent temporal patterns in information value evolution.
@@ 3.3.1 The Value of Analyst Information in the Cross-section of Stocks
Analyst information value may vary in the cross-section of stocks. Prior literature suggests that information processing and dissemination costs vary with firm size (Zeghal, 1984), and analyst opinions may trigger larger market reactions for smaller stocks due to their relatively limited public information environment (Stickel, 1995). Figure A9 illustrates this relationship through a scatter plot of information value against stock market capitalization. While the plot reveals a positive association between market equity and information value, the relationship exhibits notable nonlinearities. To formally examine this relationship, I estimate regressions of information value on stock size, incorporating firm and year fixed effects and the book-to-market ratio control.
Table 10 shows that stocks with larger market equity have significantly higher information value from analysts. The economic magnitude is substantial: estimates from column (4) indicate that a 1% increase in market equity corresponds to a 0.864% increase in analyst information value. Columns (5)-(8) reveal that textual information value exhibits greater size sensitivity compared to the combined numerical and textual information value. To better understand the mechanisms driving this positive size elasticity, I decompose the information value into two components: explained return volatility and price impact. The decomposition can be expressed logarithmically as:
@@
Panel B of Table 10 shows that higher analyst information value in larger stocks stems primarily from their greater liquidity. While analyst reports do not better explain return volatility in large stocks, the reduced price impact of liquid stocks enhances their information value.